6) A cubic crystal with a=5.60Å is set initially with the incident beam travelling along the [100] direction. The crystal is then rotated clockwise about [001] through an angle ω until the (460) reflection is first produced with Cu Kα radiation. Calculate the Bragg angle θ for the (460) reflection and determine the angle ω


1
Expert's answer
2021-03-08T03:08:33-0500

2d sin "\theta" = n "\lambda" (Braggs Law)

a = d (h2 + k2 + l2)1/2

5.60 A = d (42 + 62 + 02)1/2

d = 0.78 A

2 x 0.78 x 10-10 m x sin "\theta" = 154.18 x 10-12 m

sin "\theta" = 0.9883

"\theta" = 81°